Exercise 3.8 Constants A remote-controlled car is moving in a vacant parking lot. The velocity of the car as a function of time is given by 0 15.00 m/s-(0.0180 m/s )li+ 2.00 m/s + (0.60 m/tj

media%2Fff1%2Fff118044-1f3a-47f9-841b-69
Answer #1

t- 6.5sSec 41 V = (5-0.018 x 6.5s ,ナC2 + 0.55x 658 = 4.22)it 5.619J tu5-69 4.22 tar 0M+53.